Show that the set of rational numbers with the subspace topology of R is disconnected
1
Expert's answer
2020-05-12T18:44:42-0400

Let a"\in" R be an irrational number.

Then a"\notin" Q (rationals)

Let us suppose the sets :

P = (- "\infty" , a ) "\cap" Q

T = ( a, "\infty" ) "\cap" Q


Let x "\in P"

Let B("\in" ,x) be an open ball in of x in Q

Then, for all x in P, there exists "\in" from R such that B("\in" ,x) lies in P if "\in = a-x"


Similarly, for all x in T , there exists "\in" from R such that B ("\in" , x ) lies in T if "\in = x-a"


So , there open neighborhoods of P and T in Q, hence P and T are open sets in Q.

Now,

P "\cup" T = Q , P"\cap" T = "\varnothing" where P and T are non-empty open sets.

So, P and T are a separation in Q.

Hence, the result.